### v1

@@ Abstract
This paper considers the identification of dynamic treatment effects with panel data, in complex designs where the treatment may not be binary and may not be absorbing. We first show that under no-anticipation and parallel-trends assumptions, we can identify event-study effects comparing outcomes under the actual treatment path and under the status-quo path where all units would have kept their period-one treatment throughout the panel. Those effects can be helpful to evaluate ex-post the policies that effectively took place, and once properly normalized they estimate weighted averages of marginal effects of the current and lagged treatments on the outcome. Yet, they may still be hard to interpret, and they cannot be used to evaluate the effects of other policies than the ones that were conducted. To make progress, we impose another restriction, namely a random coefficients distributed-lag linear model, where effects remain constant over time. Under this model, the usual distributed-lag two-way-fixed-effects regression may be misleading. Instead, we show that this random coefficients model can be estimated simply. We illustrate our findings by revisiting Gentzkow et al. (2011).
@@ 1 Introduction
Most recent work on difference-in-differences have focused on simple designs, namely classical designs where some units receive a binary treatment at the same time whereas other do not, or staggered adoption designs, where units begin to receive an absorbing binary treatment at different points in time. However, more complex designs are actually prevalent. In previous work (de Chaisemartin and D’Haultfœuille, 2025), we conducted a survey of the 100 papers with the most Google Scholar citations published by the American Economic Review from 2015 to 2019. Of those, 26 use a two-way fixed effects (TWFE) regression to estimate the effect of a treatment on an outcome. [^1] We followed the long-held belief that the TWFE regression is the uncontroversial treatment-effect estimator in a DID research design. Of these 26 papers, only two have a classical design and four have an absorbing and binary treatment with variation in treatment timing. The remaining 20 have a more complex design, with a non-binary and/or non-absorbing treatment.
@@
Under our parallel-trends condition, to identify the effect of having departed from one’s status-quo treatment $\ell$ periods ago, we can form a difference-in-difference estimand comparing units that have first switched treatment $\ell$ periods ago with units that have not switched yet and that had the same treatment at period one. The corresponding event-study effects may be more challenging to interpret than in staggered adoption designs. Nonetheless, we show that they can be useful to do an ex-post cost-benefit analysis of the policies that effectively took place over the study period. Moreover, once properly normalized they estimate weighted averages of marginal effects of the current and lagged treatments on the outcome. Finally, they can be used to test relevant null hypotheses, such as whether lagged treatments affect or not the outcome. The corresponding estimators are computed by the did_multiplegt_dyn Stata and R commands.
@@
On the other hand, these event-study effects cannot be used to evaluate alternative policies, or determine the optimal policy. To make progress on these important questions, in the second part of the paper we consider a further restriction, in the form of a random coefficients distributed-lag linear model. This model is restrictive, but allows for heterogeneity in treatment effects that can be correlated with the treatment path. We first decompose the usual distributed-lag TWFE regressions under this model, and show that they usually fail to identify convex combinations of unit-specific effects in general. Then, we show that expectations of the random coefficients can be identified and estimated simply and without tuning parameters if the treatment variable is finitely supported.
@@
Finally, we illustrate our theoretical results by studying the effect of newspapers on electoral turnout in the US between 1868 and 1928, revisiting Gentzkow et al. (2011). We extend their analysis by considering a dynamic rather than static setup, thus allowing the number of past newspapers to potentially affect current turnout. Our event-study effects, relying only on no-anticipation and parallel trends assumptions, show a large positive effect of newspapers on turnout. They also suggest that the current number of newspapers has a larger effect than the lagged number of newspapers, though these parameters cannot bring a definitive answer to this question. The usual distributed-lag TWFE regression actually suggests the opposite: the coefficient on lagged newspapers is positive and significant, while the coefficient on current newspapers is small and insignificant. Yet, our decomposition of those coefficients shows that they may be unreliable if effects vary across counties. Even though somewhat imprecise, our estimates of the random coefficients distributed lag model seem to confirm the findings based on our event-study effects.
This paper contributes to a rapidly growing literature on difference-in-differences. The first part summarizes and extends results from de Chaisemartin and D’Haultfœuille (2025), a paper that extends results in Callaway and Sant’Anna (2021), Sun and Abraham (2021) and Borusyak et al. (2024) to non-binary and/or non-absorbing designs. The decomposition of the distributed-lag TWFE regression estimator is related to but different from that in de Chaisemartin and D’Haultfœuille (2025): unlike in that paper, here we assume homogeneous treatment effects over time, a natural assumption if one wants to use past policy changes to do ex-ante policy evaluation. Even under that strong additional assumption, distributed-lag TWFE regressions may not estimate convex combinations of effects. Our analysis of the random coefficient model is related to the seminal work of Chamberlain (1992), and to the more recent papers of Graham and Powell (2012) and Sasaki and Ura (2021). Compared to those more recent papers, we show that if treatment effects do not vary over time and the treatment variable is finitely supported, we obtain regular identification of some average treatment effects, which implies that these parameters can be estimated at the standard rate of convergence, without any tuning parameters.
@@
The paper is organized as follows. Section 2 presents the setup, our main assumptions and the type of designs we consider. Section 3 considers identification and estimation under no-anticipation and parallel-trends assumptions. Section 4 considers identification and estimation when one also imposes a random coefficients distributed-lag linear model. Section 5 applies the results of Sections 3 and 4 to the effect of newspapers on electoral turnout. Proofs are relegated to the appendix.
@@ 2.1 Treatment and potential outcomes
We consider a panel of $G$ groups observed at $T$ periods, respectively indexed by $g$ and $t$. Groups can be locations, like states, counties, or municipalities, but could also just be individuals or firms. We seek to identify (dynamic) effects of a treatment on a given outcome. Let $D_{g,t}$ denote the treatment of group $g$ at period $t$, with support denoted by $\mathcal{D}_{t}$, supposed to be independent of $g$. Also, let $\bm{D}_{g}=(D_{g,1},...,D_{g,T})$ denote the treatment path of group $g$, and let $\mathcal{D}$ denote its support, also assumed to be independent of $g$. We let $F_{g}$ be the first date at which $g$ switches treatment: $F_{g}=\min\{t:D_{g,t}\neq D_{g,1}\}$. If $g$ never switches, we let $F_{g}=T+1$.
For all $(d_{1},...,d_{T})\in\mathcal{D}$, let $Y_{g,t}(d_{1},...,d_{T})$ be the potential outcome of group $g$ at period $t$ if $g$ has the treatment path $(d_{1},...,d_{T})$. This potential outcome model, introduced by Robins (1986), allows for dynamic effects of lagged treatments on the current outcome, and for anticipation effects of future treatments on the current outcome. However, following most of the literature, we rule out anticipation effects hereafter:
@@ 2.2 Complex designs
(Designs with some stayers) $\mathcal{D}^{\text{r}}_{1}:=\{d\in\mathcal{D}_{1}:\;V(F|D_{1}=d)>0\}\neq\emptyset$.
We thus require that there exists an initial treatment value for which there is heterogeneity in the date at which groups change treatment for the first time. This requirement is natural in difference-in-differences contexts, and there are many applications where it holds. Still, it fails in designs without stayers, where $D_{2}\neq D_{1}$ and $F=2$ almost surely, as will for instance be the case if $D_{g,t}$ is the amount of rainfall or the average temperature in location $g$ and year $t$: all locations will experience different precipitations or temperatures in years one and two. This also fails if groups all change treatment for the first time at the same date $t_{0}$, for instance due to a universal policy affecting them all: $F=t_{0}$ almost surely. In such cases, if all groups are untreated at period one and receive heterogeneous treatment doses at $t_{0}$, the design is actually an heterogeneous adoption design and one can then use the estimators considered by de Chaisemartin et al. (2024).
@@ 2.3 Parallel trends
Assumptions 4 or 5 may be both seen as strict exogeneity conditions. One may worry that non-obsorbing designs arise because those that receive the treatment self-select in and out of it, and self-selection makes such strict exogeneity conditions implausible. However, in most of the aforementioned examples, the multiple treatment changes come from laws that are changed several times or repealed after having been enacted. Therefore, while it is important to document the reasons that led the legislator to further or cancel an initial policy change, parallel-trends assumptions are not by construction less plausible in non-absorbing designs.
@@ 3.1.1 Parameters of interest
In words, $\mathcal{D}^{\text{nc}}_{t}$ is the set of treatment paths $(d_{1},...,d_{T})$ without “crossing” until $t$: there cannot be treatment values $d_{s}$ and $d_{s^{\prime}}$ ($s,s^{\prime}\leq t$) both above and below the inital treatment value ($d_{s}<d_{1}<d_{s^{\prime}}$). Finally, let $\overline{T}_{d}:=\max\text{Supp}(F-1|D_{1}=d)$ and $\overline{T}:=\overline{T}_{D_{1}}$. The random variable $\overline{T}$ represents, for a given group, the last date at which we can find with a positive probability a “control group” with the same initial treatment value and whose treatment has not switched yet.
Now, for any $\ell\in\{1,...,T-1\}$, let
$\text{AVSQ}_{\ell}=E\left[S\left(Y_{F-1+\ell}-Y_{F-1+\ell}(D_{1},...,D_{1})\right)|D_{1}\in\mathcal{D}^{\text{r}}_{1},F-1+\ell\leq\overline{T},\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}\right].$
To interpret $\text{AVSQ}_{\ell}$, let us first suppose that $S=1$ and $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ almost surely (a.s.). Then, $\text{AVSQ}_{\ell}$ is an average expected difference between groups’ actual outcome and their counterfactual “status quo” outcome if their treatment had always remained equal to their period-one value $D_{1}$, $\ell$ periods after the first switch occurs. Because of this, we refer to $\text{AVSQ}_{\ell}$ as an actual-versus-status-quo (AVSQ) event-study effect. The expectation is over groups for which $D_{1}\in\mathcal{D}^{\text{r}}_{1},F-1+\ell\leq\overline{T}$. Intuitively, and as shown below, these are the groups for which the expected counterfactual outcome can be identified.
In a binary and staggered adoption design, $S=1$ and $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ a.s., $\mathcal{D}^{\text{r}}_{1}=\{0\}$, and thus, $\text{AVSQ}_{\ell}$ boils down to
@@
The interpretation of $\text{AVSQ}_{\ell}$ is more delicate in non-binary and/or non-staggered designs, even if $S=1$ and $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ a.s. For instance, if Eq. (1) holds, so that $D_{t}=1\{E\geq t\geq F\}$ for some random $E$ and $F$, $\text{AVSQ}_{\ell}$ is a weighted average of
@@
We introduce the condition $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ in the definition of $\text{AVSQ}_{\ell}$ to ensure that it satisfies the following “no-sign reversal” property (Imbens and Angrist, 1994; Small et al., 2017):
If $(d_{1},...,d_{t})\geq(d^{\prime}_{1},...,d^{\prime}_{t})$ (where the inequality should be understood componentwise) $\Rightarrow Y_{t}(d_{1},...,d_{t})\geq Y_{t}(d^{\prime}_{1},...,d^{\prime}_{t})$ a.s., then $\text{AVSQ}_{\ell}\geq 0$.
@@
$=$ $E\left[Y_{3}(1,2,1)-Y_{3}(1,1,1)|F=2,D_{1}=1\right]-E\left[Y_{3}(1,2,1)-Y_{3}(1,2,0)|F=2\right].$
@@ 3.1.2 Identification and estimation
$\text{AVSQ}_{\ell}=$ $E\left[S(Y_{F-1+\ell}-Y_{F-1}-\Delta_{\ell}(D_{1},F))|\bm{D}\in\mathcal{D}^{\ell}\right],$ (4)
with $\Delta_{\ell}(d_{1},f)=E[Y_{f-1+\ell}-Y_{f-1}|D_{1}=d,F>f-1+\ell]$. Moreover, $P(\bm{D}\in\mathcal{D}^{\ell})>0$ holds for at least $\ell=1$.
@@
When $\mathcal{S}_{\ell}=\emptyset$, we simply let $\widehat{\text{AVSQ}}_{\ell}=0$. The set $\mathcal{C}_{g,\ell}$ includes the congrol groups for $g$, namely the groups with the same period-one treatment as switcher $g$ and whose treatment has not changed yet at period $F_{g}-1+\ell$. Hence, $\mathcal{S}_{\ell}$ denotes the subset of switchers $g$ satisfying the no-crossing condition and for which a control group can be found. The condition $g\in\mathcal{S}_{\ell}$ can therefore be seen as the finite-sample counterpart of $\bm{D}_{g}\in\mathcal{D}^{\ell}$.
By slightly adapting the proof of Theorem 1 in (de Chaisemartin and D’Haultfœuille, 2025), we can prove that $\widehat{\text{AVSQ}}_{\ell}$ is consistent and asymptotically normal for $\text{AVSQ}_{\ell}$, as $G$ tends to infinity, under mild restrictions (de Chaisemartin and D’Haultfœuille, 2025).

### v3

@@ Abstract
This paper considers the identification of dynamic treatment effects with panel data, in complex designs where the treatment may not be binary and may not be absorbing. We first show that under no-anticipation and parallel-trends assumptions, we can identify event-study effects comparing outcomes under the actual treatment path and under the status-quo path where all units would have kept their period-one treatment throughout the panel. Those effects can be helpful to evaluate ex-post the policies that effectively took place, and once properly normalized they estimate weighted averages of marginal effects of the current and lagged treatments on the outcome. Yet, they may still be hard to interpret, and they cannot be used to evaluate the effects of other policies than the ones that were conducted. To make progress, we impose another restriction, namely a random coefficients distributed-lag linear model, where effects remain constant over time. Under this model, the usual distributed-lag two-way-fixed-effects regression may be misleading. Instead, we show that this random coefficients model can be estimated simply. We illustrate our findings by revisiting Gentzkow, Shapiro, and Sinkinson (2011).
@@ 1 Introduction
Most work on difference-in-differences has focused on simple designs, namely classical designs where some units receive a binary treatment at the same time whereas other do not, or staggered adoption designs, where units begin to receive an absorbing binary treatment at different points in time, or heterogeneous adoption designs where all units are initially untreated before the treated units start receiving heterogeneous doses. However, more complex designs are actually prevalent. In previous work (de Chaisemartin and D’Haultfœuille, 2025), we conducted a survey of the 100 papers with the most Google Scholar citations published by the American Economic Review from 2015 to 2019. Of those, 26 use a two-way fixed effects (TWFE) regression to estimate the effect of a treatment on an outcome. [^1] Counting TWFE regressions is in line with the long-held belief that this regression is the uncontroversial treatment-effect estimator in a DID research design. Of these 26 papers, only two have a classical design, four have an absorbing and binary treatment with variation in treatment timing, and two have an heterogeneous adoption design. The remaining 18 have a more complex design, with a non-binary and/or non-absorbing treatment, and where units may receive heterogeneous treatment doses even at period one.
@@
Under our parallel-trends condition, to identify the effect of having departed from one’s status-quo treatment $\ell$ periods ago, we can form a difference-in-difference estimand comparing units that have first switched treatment $\ell$ periods ago with units that have not switched yet and that had the same treatment at period one. The corresponding event-study effects may be more challenging to interpret than in staggered adoption designs. Nonetheless, we show that they can be useful to do an ex-post cost-benefit analysis of the policies that effectively took place over the study period. Moreover, once properly normalized they estimate weighted averages of marginal effects of the current and lagged treatments on the outcome. Finally, they can be used to test relevant null hypotheses, such as whether lagged treatments affect the outcome or not. The corresponding estimators are computed by the did_multiplegt_dyn Stata, R, and Python commands.
@@
Still, these event-study effects cannot be used to separate the marginal effects of the current treatment and of some specific treatment lags. Moreover, they cannot be used to evaluate alternative policies, or determine the optimal policy. To make progress on these important questions, we consider in the second part of the paper a further restriction, in the form of a random coefficients distributed-lag linear model. This model is restrictive, but allows for heterogeneity in treatment effects that can be correlated with the treatment path. We first decompose the usual distributed-lag TWFE regressions under this model, and show that they usually fail to identify convex combinations of unit-specific effects. Then, we show that expectations of the random coefficients can be identified and estimated simply and without tuning parameters if the treatment variable takes a finite number of values. The corresponding estimators are computed by the dist_lag_het R package, available at https://github.com/chaisemartinPackages/dist_lag_het.
@@
Finally, we illustrate our theoretical results by studying the effect of newspapers on electoral turnout in the US between 1868 and 1928, revisiting Gentzkow, Shapiro, and Sinkinson (2011). We extend their analysis by considering a dynamic rather than static setup, thus allowing the number of past newspapers to potentially affect current turnout. Our event-study effects, relying only on no-anticipation and parallel trends assumptions, show a large positive effect of newspapers on turnout. Though these parameters cannot bring a definitive answer to this question, they also suggest that the current number of newspapers has a larger effect than the lagged number of newspapers, and our estimates of the random coefficients distributed lag model confirm this finding. On the other hand, the distributed-lag TWFE regression leads to the opposite conclusion: the coefficient on lagged newspapers is positive and significant, while the coefficient on current newspapers is small and insignificant. Our decomposition of those coefficients shows that they may be unreliable if effects vary across counties.
This paper contributes to a rapidly growing literature on difference-in-differences (DID). In that literature, most papers have focused on designs with a binary and absorbing treatment (de Chaisemartin and D’Haultfœuille, 2020; Callaway and Sant’Anna, 2021; Sun and Abraham, 2021; Borusyak et al., 2024) or on heterogeneous adoption designs where all units are untreated at period one before treated units start receiving heterogeneous doses (Fricke, 2017; de Chaisemartin and D’Haultfœuille, 2020; Callaway et al., 2021; de Chaisemartin et al., 2024). The first part of the paper summarizes and extends results from de Chaisemartin and D’Haultfœuille (2025), that can accommodate complex designs with a non-binary and/or non-absorbing treatment, and where units may receive heterogeneous treatment doses even at period one. We emphasize that under a parallel-trends assumption alone, we can identify weighted averages of the effect of the current and lagged treatments on the outcome, without being able to separately estimate those effects. This motivates our analysis of the random coefficients distributed-lag model in the second part of the paper. There, we show that the work of Chamberlain (1992), Arellano and Bonhomme (2012), and Graham and Powell (2012) can fruitfully be applied to estimate distributed-lag models with heterogeneous effects. Compared to these papers, we highlight the fact that averages of the random coefficients can be estimated simply and without tuning parameters if the treatment variable takes a finite number of values. We also provide identification conditions for the average of each random coefficient, rather than the average of the whole vector.
@@
The paper is organized as follows. Section 2 presents the setup, our main assumptions and the type of designs we consider. Section 3 considers identification and estimation under no-anticipation and parallel-trends assumptions. Section 4 considers identification and estimation when one also imposes a random coefficients distributed-lag linear model. Section 5 applies the results of Sections 3 and 4 to estimate the effect of newspapers on electoral turnout. Proofs are relegated to the appendix.
@@ 2.1 Treatment and potential outcomes
We consider a panel of $G$ groups observed at $T$ periods, respectively indexed by $g$ and $t$. Groups can be locations, like states, counties, or municipalities, but could also just be individuals or firms. We seek to identify (dynamic) effects of a treatment on a given outcome. Let $D_{g,t}$ denote the treatment of group $g$ at period $t$, with support denoted by $\mathcal{D}_{t}$, supposed to be independent of $g$. Also, let $\bm{D}_{g}=(D_{g,1},...,D_{g,T})$ denote the treatment path of group $g$, and let $\mathcal{D}$ denote its support, also assumed to be independent of $g$. We let $F_{g}$ be the first date at which $g$ switches treatment: $F_{g}=\min\{t\geq 2:D_{g,t}\neq D_{g,1}\}$. If $g$ never switches, we let $F_{g}=T+1$.
For all $(d_{1},...,d_{T})\in\mathcal{D}$, let $Y_{g,t}(d_{1},...,d_{T})$ be the potential outcome of group $g$ at period $t$ if $g$ has the treatment path $(d_{1},...,d_{T})$. This potential outcome model, introduced by Robins (1986), allows for dynamic effects of lagged treatments on the current outcome, and for anticipation effects of future treatments on the current outcome. However, following most of the literature, we rule out anticipation effects hereafter: [^2] The notation $Y_{g,t}(d_{1},...,d_{T})$ also implicitly rules out dependence on treatments that may have occurred before period 1; see Section 1.8 in the web appendix of de Chaisemartin and D’Haultfœuille (2025) for a discussion on this issue.
@@ 2.2 Complex designs
(Designs with some stayers) $\mathcal{D}^{\text{r}}_{1}:=\{d\in\mathcal{D}_{1}:\;V(F|D_{1}=d)>0\}\neq\emptyset$. [^3] In case the distribution of $D_{1}$ is continuous, $V(F|D_{1}=d)$ is only defined on a set of $d$’s of measure 1. Then, we implicitly assume in what follows that there exists a continuous version of $d\mapsto V(F|D_{1}=d)$.
We thus require that there exists an initial treatment value for which there is heterogeneity in the date at which groups change treatment for the first time. This requirement is natural in difference-in-differences contexts, and there are many applications where it holds. Still, it fails in designs without stayers, where $D_{2}\neq D_{1}$ and $F=2$ almost surely, as will for instance be the case if $D_{g,t}$ is the amount of rainfall or the average temperature in location $g$ and year $t$: all locations will experience different precipitations or temperatures in years one and two. This also fails if groups all change treatment for the first time at the same date $t_{0}$, for instance due to a universal policy affecting them all: $F=t_{0}$ almost surely. In such cases, if all groups are untreated at period one and receive heterogeneous treatment doses at $t_{0}$, the design is actually an heterogeneous adoption design and one can then use the estimators considered by de Chaisemartin, Ciccia, D’Haultfœuille, and Knau (2024).
@@ 2.3 Parallel trends
Assumptions 4 or 5 may be both seen as strict exogeneity conditions. One may worry that non-absorbing designs arise because those that receive the treatment self-select in and out of it, and self-selection makes such strict exogeneity conditions implausible. However, in most of the aforementioned examples, the multiple treatment changes come from laws that are changed several times or repealed after having been enacted. Therefore, while it is important to document the reasons that led the legislator to further or cancel an initial policy change, parallel-trends assumptions are not by construction less plausible in non-absorbing designs.
@@ 3.1.1 Parameters of interest
In words, $\mathcal{D}^{\text{nc}}_{t}$ is the set of treatment paths $(d_{1},...,d_{T})$ without “crossing” until $t$: there cannot be treatment values $d_{s}$ and $d_{s^{\prime}}$ ($s,s^{\prime}\leq t$) both above and below the initial treatment value ($d_{s}<d_{1}<d_{s^{\prime}}$). Finally, let $\overline{T}_{d}:=\max\text{Supp}(F-1|D_{1}=d)$ and $\overline{T}:=\overline{T}_{D_{1}}$. The random variable $\overline{T}$ represents, for a given group, the last date at which we can find with a positive probability a “control group” with the same initial treatment value and whose treatment has not switched yet.
Now, for any $\ell\in\{1,...,T-1\}$ for which $P(D_{1}\in\mathcal{D}^{\text{r}}_{1},F-1+\ell\leq\overline{T},\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell})>0$, let
$\text{AVSQ}_{\ell}=E\left[S\times\left(Y_{F-1+\ell}-Y_{F-1+\ell}(D_{1},...,D_{1})\right)|D_{1}\in\mathcal{D}^{\text{r}}_{1},F-1+\ell\leq\overline{T},\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}\right].$
To interpret $\text{AVSQ}_{\ell}$, let us first suppose that $S\geq 0$ and $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ almost surely (a.s.). Then, $\text{AVSQ}_{\ell}$ is an average expected difference between groups’ actual outcome and their counterfactual “status quo” outcome if their treatment had always remained equal to their period-one value $D_{1}$, $\ell$ periods after the first switch occurs. Because of this, we refer to $\text{AVSQ}_{\ell}$ as an actual-versus-status-quo (AVSQ) event-study effect. The expectation is over groups for which $D_{1}\in\mathcal{D}^{\text{r}}_{1},F-1+\ell\leq\overline{T}$. Intuitively, and as shown below, these are the groups for which the expected counterfactual outcome can be identified.
In a binary and staggered adoption design, $S\geq 0$ and $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ a.s., $\mathcal{D}^{\text{r}}_{1}=\{0\}$, and thus, $\text{AVSQ}_{\ell}$ boils down to
@@
The interpretation of $\text{AVSQ}_{\ell}$ is more delicate in non-binary and/or non-staggered designs, even if $S\geq 0$ and $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ a.s. For instance, if Eq. (1) holds, so that $D_{t}=1\{E\geq t\geq F\}$ for some random $E$ and $F$, $\text{AVSQ}_{\ell}$ is a weighted average of
@@
We introduce the condition $\bm{D}\in\mathcal{D}^{\text{nc}}_{F-1+\ell}$ in the definition of $\text{AVSQ}_{\ell}$ to ensure that it satisfies the following “no-sign reversal” property (Imbens and Angrist, 1994; Small, Tan, Ramsahai, Lorch, and Brookhart, 2017):
If $(d_{1},...,d_{t})\geq(d^{\prime}_{1},...,d^{\prime}_{t})$ (where the inequality should be understood component-wise) $\Rightarrow Y_{t}(d_{1},...,d_{t})\geq Y_{t}(d^{\prime}_{1},...,d^{\prime}_{t})$ a.s., then $\text{AVSQ}_{\ell}\geq 0$.
@@
$=$ $E\left[Y_{3}(1,2,1)-Y_{3}(1,1,1)|F=2\right]-E\left[Y_{3}(1,2,1)-Y_{3}(1,2,0)|F=2\right].$
@@ 3.1.2 Identification and estimation
$\text{AVSQ}_{\ell}=$ $E\left[S\times(Y_{F-1+\ell}-Y_{F-1}-\Delta_{\ell}(D_{1},F))|\bm{D}\in\mathcal{D}^{\ell}\right],$ (4)
with $\Delta_{\ell}(d_{1},f)=E[Y_{f-1+\ell}-Y_{f-1}|D_{1}=d_{1},F>f-1+\ell]$. Finally, $P(\bm{D}\in\mathcal{D}^{\ell})>0$ holds for at least $\ell=1$.
@@
When $\mathcal{S}_{\ell}=\emptyset$, we simply let $\widehat{\text{AVSQ}}_{\ell}=0$. The set $\mathcal{C}_{g,\ell}$ includes the control groups for $g$, namely the groups with the same period-one treatment as switcher $g$ and whose treatment has not changed yet at period $F_{g}-1+\ell$. Hence, $\mathcal{S}_{\ell}$ denotes the subset of switchers $g$ satisfying the no-crossing condition and for which a control group can be found. The condition $g\in\mathcal{S}_{\ell}$ can therefore be seen as the finite-sample counterpart of $\bm{D}_{g}\in\mathcal{D}^{\ell}$.
By slightly adapting the proof of Theorem 1 in de Chaisemartin and D’Haultfœuille (2025), we can prove that $\widehat{\text{AVSQ}}_{\ell}$ is consistent and asymptotically normal for $\text{AVSQ}_{\ell}$, as $G$ tends to infinity, under mild restrictions (de Chaisemartin and D’Haultfœuille, 2025).